OpenStudy (anonymous):

Find a vector equation and parametric equations for the line segment that joins P(2, 0, 0) Q(6, 2, 2)

OpenStudy (anonymous):

Remember that to get vector equation you need any point that lies on to the line and a vector that goes in the same direction.

OpenStudy (anonymous):

You could pick point P or point Q.

OpenStudy (across):

You could try\[\mathbf{r}(t)=\mathbf{P}+t(\mathbf{Q}-\mathbf{P}),\qquad0\leq t\leq1.\]Notice that if \(t=0\), then \(\mathbf{r}(t)=\mathbf{P}\), and if \(t=1\), then \(\mathbf{r}(t)=\mathbf{Q}\).

OpenStudy (anonymous):

How do you make the letters bold @across ?

OpenStudy (anonymous):

@across Do I just plug in the values P and Q for that formula?

OpenStudy (across):

\mathbf{}

OpenStudy (anonymous):

just write Large before you type

OpenStudy (anonymous):

\[\vec{r}=\large{P}\]

OpenStudy (anonymous):

@zordoloom did you get it now ?

OpenStudy (anonymous):

So the formula that across provided, Can I use that to find the equations?

OpenStudy (anonymous):

first find the required parallel vector to the line PQ=<6-2,2-0,2-0)= <4,2,2>

OpenStudy (anonymous):

yes you can use that .

OpenStudy (anonymous):

so, i end up with r(t)=(2,0,0)+t(4,2,-2). Is this correct?

OpenStudy (anonymous):

r(t)=(2,0,0)+t(4,2,2 ) not -2 !!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

welcome :)